OpenStudy (anonymous):

(x)= -80e^(-.1x)-40x and it says to determind relative max or min or inflection points if they exist using the second derivative technique

OpenStudy (anonymous):

Think of it this way, if you have 15 (x+2)^4 = 0 you can divide both sides by 15 to give you (x+2)^4 = 0 then you can think of (x+2)^4 = 0 as (x+2)(x+2)(x+2)(x+2)=0 the only way you can get that whole thing to be 0 is if (x+2) = 0 which means that x = -2 I hope that makes sense

OpenStudy (anonymous):

For this question, try to think of something that you can plug into "X" that will make (x+2) equal to 0. You can ignore the 15 as well as the ^4 because as long as (x+2)=0, the rest will be 0 too! Ex: 15(0)^4 will always equal 0 :)

OpenStudy (anonymous):

Ok what is the inflection point or relative max or min? Thanks again

OpenStudy (anonymous):

this is for pre-calculus

OpenStudy (anonymous):

do you know how to take the derivative?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

basically i need to determine any relative minimum /maximum or inflection points if they exist using the second derivative technique for this equation f(x)=3(x+2)^5-2

OpenStudy (anonymous):

ok so you need the derivative and set it to be 0 then solve for x to find the x values of the relative max and min. The derivative will be 0 at these points because the tangent will be horizontal and have a slope of 0.

OpenStudy (anonymous):

basically i need to determine any relative minimum /maximum or inflection points if they exist using the second derivative technique for this equation f(x)=3(x+2)^5-2

OpenStudy (anonymous):

yeah to find the relative max and min you need the first derivative and set it to be 0 for the inflection points you need the second derivative

OpenStudy (anonymous):

f`(x)= 15(x+2)^4 is the first derivative

OpenStudy (anonymous):

now setting that equals to zero, you said x= -2

OpenStudy (anonymous):

and the second derivative is f``(x)= 60(x+2)^3

OpenStudy (anonymous):

now what ?

OpenStudy (anonymous):

setting that to be 0 should give you the inflection point. which in this case is x = -2 as well

OpenStudy (anonymous):

so what is the inflection point or relative max or min?

OpenStudy (anonymous):

inflection points*

OpenStudy (anonymous):

well you have the x valve. All you need to do is sub it into the original function to get your y value. (-2,-2) would be your relative max/min as well as your inflection point

OpenStudy (anonymous):

if you look at the graph for the function you can see that this is the case since (-2,-2) is the only point where the inflection changes and can also be classified as a relative max/min. http://www.wolframalpha.com/input/?i=y%3D3%28x%2B2%29^5-2

OpenStudy (anonymous):

okay one more question for a different problem. What if the first derivative is f`(x)= 8 Then what is the relative max or max or inflection points?

OpenStudy (anonymous):

The problem states this: f(x)= -x^4/4 + 9x ^2/2 and it says to determine any relative max or min or inflection pints if they exist using the second derivative technique.

OpenStudy (anonymous):

could you put brackets in that? I can't understand what goes where

OpenStudy (anonymous):

f(x)= -x^(4/4) + 9x^(2/2)

OpenStudy (anonymous):

the first derivative to that equation is f`(x)= 8

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

f(x)= -x^(4/4) + 9x^(2/2) is basically -x + 9x or just 8x so the derivative would be 8. If you think about the line y=8x then it becomes clear that there are no relative max/min or inflection points since the line is straight. When you try and do the derivative test you just get 0=8 which is impossible to satisfy.

OpenStudy (anonymous):

okay yeah its a straight line. Thanks man soo much. And this is the last question : f(x)= -80e^(-.1x)-40x and it says to determind relative max or min or inflection points if they exist using the second derivative technique

OpenStudy (anonymous):

ok so from the graph we can see that there should be a relative max, but not inflection points (since it's always concave down) http://www.wolframalpha.com/input/?i=f%28x%29%3D+-80e^%28-.1x%29-40x

OpenStudy (anonymous):

the first derivative to that one that i found was i think (not sure correct me if I'm wrong) f`(x)=8e^(-0.1x)-40

OpenStudy (anonymous):

and the second derivative was i think : f``(x)= -0.8e^(-0.1x)

OpenStudy (anonymous):

So what is the relative max points?

OpenStudy (anonymous):

well set the first derivative to be 0 and solve for x

OpenStudy (anonymous):

yup and that shows x=-16.09

OpenStudy (anonymous):

so what did you get as the relative max pt

OpenStudy (anonymous):

yeah you should have got x = -10log(5) which is -16.09

OpenStudy (anonymous):

so what is the inflection point or the relative max or minimum?

OpenStudy (anonymous):

well that value right there is the x value for the relative max of that function. To get the y you need to sub it back into f(x) like before.

OpenStudy (anonymous):

what did you get as the y value

OpenStudy (anonymous):

for the inflection point you have 0 = -0.8e^(-0.1x) but it can never be 0 since -0.8e^(-0.1x) asymptotically approaches 0 (but never reaches it)

OpenStudy (anonymous):

so is it an inflection pt or relative max or min?

OpenStudy (anonymous):

and what is the y-value?

OpenStudy (anonymous):

well the inflection point doesn't exist since -0.8e^(-0.1x) can never be 0 so there are no points for the second derivative test

OpenStudy (anonymous):

i know x is the -16.09 but what is the y-value you got

OpenStudy (anonymous):

so would it be a relative min or the max?

OpenStudy (anonymous):

the y value is 243.775 and it is a max.

OpenStudy (anonymous):

but keep in mind that the first derivative test doesn't tell you that is is a max or a min. Just that it is possibly a relative max or min

OpenStudy (anonymous):

I know just from looking at the graph. Otherwise you would need to test some values if I remember correctly

OpenStudy (anonymous):

I got max as well. but i got -1,043.42 as the y value??

OpenStudy (anonymous):

f(-10log(5)) = (-80e^(-.1*(-10log(5))) - 40*-10log(5)) are you sure you are doing this?

OpenStudy (anonymous):

no i did it someway else

OpenStudy (anonymous):

i did ln

OpenStudy (anonymous):

but im sure yours is right

OpenStudy (anonymous):

thanks again man I really appreciate it! And hopefully I get a 100% on this take home quiz homework

OpenStudy (anonymous):

You've helped me A LOT!

OpenStudy (anonymous):

hopefully everything I did 100% correct haha. It's been some time since I've done calculus. I'm glad I could help though!